OpenStudy (anonymous):

How many integers between 10 and 60, inclusive, can be evenly divided by neither 2 nor 3?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

total integers =?

OpenStudy (anonymous):

51 integers

hartnn (hartnn):

of them no. of integers divided by 2=?

OpenStudy (anonymous):

25

hartnn (hartnn):

and no. of integers divided by 3=?

hartnn (hartnn):

17?

OpenStudy (anonymous):

yes 17

hartnn (hartnn):

16 i guess

hartnn (hartnn):

12 is first, 60 is last ....

OpenStudy (anonymous):

yes it will be 17

hartnn (hartnn):

ok, 17....and no. of integers divisible by both 2 and 3 =?? that is no. of integers divisible by 6=?

OpenStudy (ajprincess):

60=12+(n-1)3 60-12=3n-3 48+3=3n 51=3n n=17 60=10+(n-1)2 50=2n-2 52=2n 26=n

OpenStudy (anonymous):

just a minute the number of integers divided by 2 is 26

hartnn (hartnn):

yes,like @ajprincess explained

hartnn (hartnn):

what about 6?

hartnn (hartnn):

9?>

OpenStudy (anonymous):

is 34 the final answer?

hartnn (hartnn):

51-17-26-9=?

hartnn (hartnn):

sorry +9

hartnn (hartnn):

i m getting 17

OpenStudy (anonymous):

ok.. thanks

hartnn (hartnn):

welcome ^_^